OpenStudy (anonymous):

Simplify (2x + 3)(x – 4) = ? Will give medal. P.s. I want a full explanation as to how you solved this problem. Please and Thanks a Ton! ^_^

OpenStudy (aravindg):

You know how to open up (a+b)(c+d)?

OpenStudy (aravindg):

(a+b)(c+d)=ac+ad+bc+bd

OpenStudy (anonymous):

i want a full explenation as to how to solve it, including opening it up. Please and thanks. ^_^

OpenStudy (aravindg):

Don't expect us to give out a full worked out answer, I can give you inputs to help you get to the answer.

OpenStudy (anonymous):

i dont want the answer. ^_^ i just want how to solve it. i can do the rest

OpenStudy (anonymous):

@AravindG

OpenStudy (aravindg):

Use the formula I showed you . (a+b)(c+d)=ac+ad+bc+bd example (3x+4)(x+5)=3x*x+3x*5+4*x+4*5

OpenStudy (anonymous):

ok, thanks. ^_^

OpenStudy (aravindg):

Did you get it?

OpenStudy (anonymous):

yerp, im solving it right now ^_^

OpenStudy (anonymous):

@AravindG is it 2x^2 + 5x – 12 ?

OpenStudy (aravindg):

+5x?

OpenStudy (anonymous):

or is it 2x^2 – 5x – 12 ?

OpenStudy (anonymous):

oh, yah, it's -5x isnt it?

OpenStudy (aravindg):

Yep -8x+3x=-5x

OpenStudy (anonymous):

kk, thanks! ^_^